A Bibliography of Publications in Biometrics: 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 September 2024 Version 1.02

Title word cross-reference

2 × 2 [13]. p [93, 47]. p^{3} [93]. p^{n} [47, 55]. p^{s} [55].

1946 [7].

Abstracts [11, 20, 39, 59, 87]. Activity [53]. Additivity [105]. Allergens [54]. Allocation [102]. Allocution [37]. Allometry [46]. Alyce [71]. Analyses [71]. Analysis [95, 62, 2, 12, 102, 92, 72, 3, 93, 90]. Annual [5, 50]. Antibiotic [115]. Application [102]. Approximation [106]. Assay [122, 112, 114]. Assaying [115]. Assays [13, 116, 113, 56]. Association [23]. Assumptions [2, 3]. August [7]. Australia [33]. Average [75].

Based [95]. between [12, 94]. Binomial [95, 94]. Bio [122, 112]. Bio-Assay [122, 112]. Bioassay [92, 64]. Biological [13, 29, 114, 116, 113, 30, 46]. Biologique [46]. Biometric [33, 34, 31, 49, 60, 79, 100, 110, 119, 38, 67, 36, 88]. Biometrics [37, 5, 50]. Biometrie [37]. Biparental [75]. block [32]. Blocks [47, 55, 93]. Blood [56]. both [103].

California [51]. Capacity [74]. Catch [83, 104]. Cattle [54]. Changes [104]. Chemical [102, 71]. Chiasma [73, 45]. Choice [122, 112]. Closing [37]. Cloture [37]. Clover [71]. Committee [6]. Comparing [90]. Comparison [54]. Comparisons [72, 22]. Complement [64]. Complement-Fixation [64]. Components [75, 102]. Computation [116]. Conference [37, 37]. consequences [3]. Constitution [36]. Contrasted [24]. Correction [28, 7, 122, 123]. Correlation [23]. Cotton [63]. Council [38]. Counts [95, 14]. Covariance [62, 12, 72]. Cranial [74]. Curves [64].

D’Allometrie [46]. Data [92]. Date [63]. Day [56]. Degree [75, 105]. Denmark [31]. Design [63, 54]. Designs [47, 55, 32, 93, 55]. Development [21]. Developments [34, 31]. Diffusion [115]. Dimensional [115]. Discussion [35]. Distribution [95]. Dominance [75].

Editorial [6]. Egg [70]. Eggs [14]. England [104]. Errata [73]. error [103]. Escapements [91]. Estimate [63]. Estimating [75, 70, 91]. Estimation [29]. Experimental [63, 54]. Experiments [13]. Extinction [92].

Factorial [13]. First [37]. Fishery [84]. Fitting [103]. Fixation [64]. Formation [73, 45]. Formula [74]. Formulas [115]. Fourfold [23]. Freedom [105]. French [37, 46]. Front [44, 52, 61, 69, 81, 89, 101, 111, 121]. Fundamental [115].

General [47, 55, 93]. Genetic [73, 75, 45]. Groups [13].

Haddock [104]. Health [23]. Hospital [21].

II [55]. III [93]. Illustration [64]. Incomplete [13, 32]. India [32]. individual [90, 22]. Infection [106]. Information [43, 80, 120]. Insect [95]. Insulin [56]. International [37]. Internationale [37]. Interval [106]. Introduction [47]. Involving [116].

Laboratory [30]. Lake [84]. Lakes [83]. Lattice [47, 55, 93]. Lattices [82, 1]. Life [24]. line [103]. logarithmic [94].

Marine [30]. Matter [44, 52, 61, 69, 81, 89, 101, 111, 121]. Meaning [113]. means [90]. Meeting [5, 50]. Meetings [38]. Metameter [122, 112]. Method [12, 115]. Methods [91, 22].

Negative [95, 94]. Nervous [53]. Nested [102]. Netherlands [34]. News [42, 68]. No [7]. Non [72, 105]. Non-Additivity [105]. Non-Orthogonal [72]. Nonparallel [64]. Note [6]. Notes [42, 68].

Observed [24]. One [115, 105, 56]. One-Dimensional [115]. Optimum [63, 102, 83]. Organization [53]. Orthogonal [72].

Parallel [64]. Part [12]. Partial [23]. Planting [63]. Plots [47, 55, 93]. Poisson [94]. Population [104]. Populations [75, 29]. Poultry [70]. Power [47, 55, 93]. Practices [21]. Prediction [74]. Preliminary [1]. Première [37]. Prime [47, 55, 93]. Prime-Power [47, 55, 93]. Probability [22]. Problems [23, 83]. Production [70]. Progenies [75]. Public [23].

Quantitative [73, 116, 64, 45]. Query [28, 123, 15, 85, 7, 66, 96, 97, 98, 107, 108, 58, 26, 99, 10, 8, 16, 41, 117, 76, 17, 40, 9, 86, 19, 18, 25, 27, 48, 57, 65, 77, 78, 109, 118].

Rabbit [56]. Random [70]. ranking [22]. Recombination [73, 45]. Rectangular [82, 1]. Relates [23]. Relation [12, 46, 94]. Relationship [104, 46]. Remarks [37]. replicates [93]. Report [38, 1]. Response [122, 112, 116]. Results [113]. Routine [116]. Runs [91].

Sa [46]. Salmon [91]. Sample [56]. Sampling [102, 70, 71]. satisfied [3]. Section [5, 50]. series [94]. Sessions [51]. Significance [46]. Signification [46]. Size [104]. Small [83]. Society [49, 60, 79, 100, 110, 119, 38, 67, 36, 88]. Some [3]. Specific [24]. Squares [55]. Statistical [51, 95, 106, 21, 14, 53, 46]. Statistics [84]. Statistique [46]. straight [103]. subject [103]. Substances [115]. Summer [51]. Survivorship [24]. Systematic [70]. Systems [64].

Table [23]. Tables [24, 22]. Test [64, 56]. Their [75]. Theory [73, 47, 55, 93, 45]. Time [92, 24]. Time-Specific [24]. Total [91]. Transformations [4]. Transformed [64]. Treatment [14, 23]. Trichostrongylid [14]. Triple [82]. Trout [84]. Tuberculins [114].

Underlying [2]. University [51]. Use [4, 13, 75, 21, 64]. Used [63].

Validity [113]. variables [103]. Variance [75, 2, 102, 3, 90]. Varieties [47, 55, 93]. Volume [43, 80, 120].

Welcome [30]. Whitefish [83]. Whole [12]. Work [33, 21, 32].

References

Harshbarger:1946:PRR

[1] Boyd Harshbarger. Preliminary report on the rectangular lattices. Biometrics, 2(??):115–119, 1946. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eisenhart:1947:AUA

[2] Churchill Eisenhart. The assumptions underlying the analysis of variance. Biometrics, 3(1):1–21, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001534.

Cochran:1947:SCW

[3] W. G. Cochran. Some consequences when the assumptions for the analysis of variance are not satisfied. Biometrics, 3(1):22–38, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001535.

Bartlett:1947:UT

[4] M. S. Bartlett. The use of transformations. Biometrics, 3(1):39–52, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001536.

Anonymous:1947:AMB

[5] Anonymous. Annual meeting of the Biometrics Section. Biometrics, 3(1): 53, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001537.

Anonymous:1947:NEC

[6] Anonymous. A note from the Editorial Committee. Biometrics, 3(1): 53, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001538.

Anonymous:1947:CQN

[7] Anonymous. Correction in Query No. 33, August, 1946. Biometrics, 3(1): 54, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001539.

Crump:1947:Q

[8] S. Lee Crump. 43. Query. Biometrics, 3(1):54, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001540.

Lush:1947:Q

[9] Jay L. Lush. 44. Query. Biometrics, 3(1):54–55, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001541.

Brown:1947:Q

[10] G. W. Brown. 45. Query. Biometrics, 3(1):55–57, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001542.

Anonymous:1947:Aa

[11] Anonymous. Abstracts. Biometrics, 3(1):58–62, March 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001543.

Fisher:1947:ACM

[12] R. A. Fisher. The analysis of covariance method for the relation between a part and the whole. Biometrics, 3(2):65–68, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001641.

Bliss:1947:FEI

[13] C. I. Bliss. 2 × 2 factorial experiments in incomplete groups for use in biological assays. Biometrics, 3(2):69–88, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001642.

Emik:1947:STC

[14] L. Otis Emik. Statistical treatment of counts of trichostrongylid eggs. Biometrics, 3(2):89–93, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001643.

Anderson:1947:Q

[15] R. L. Anderson. 46. Query. Biometrics, 3(2):94–97, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001644. See correction [28].

Darroch:1947:Q

[16] J. G. Darroch. 47. Query. Biometrics, 3(2):97–98, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001645.

Kempthorne:1947:Qa

[17] Oscar Kempthorne. 48. Query. Biometrics, 3(2):98–99, June 1947. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001646.

Snedecor:1947:Qa

[18] George W. Snedecor. 49. Query. Biometrics, 3(2):99–101, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001647.

Schumacher:1947:Q

[19] F. X. Schumacher. 50. Query. Biometrics, 3(2):101–103, June 1947. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001648.

Anonymous:1947:Ab

[20] Anonymous. Abstracts. Biometrics, 3(2):104–105, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001649.

Densen:1947:DUS

[21] Paul M. Densen. The development and use of statistical practices in hospital work. Biometrics, 3(3):109–118, September 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001945.

Wilcoxon:1947:PTI

[22] Frank Wilcoxon. Probability tables for individual comparisons by ranking methods. Biometrics, 3(3):119–122, September 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001946.

Lombard:1947:TFT

[23] Herbert L. Lombard and Carl R. Doering. Treatment of the fourfold table by partial association and partial correlation as it relates to public health problems. Biometrics, 3(3):123–128, September 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001947.

Merrell:1947:TSL

[24] Margaret Merrell. Time-specific life tables contrasted with observed survivorship. Biometrics, 3(3):129–136, September 1947. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/3001948.

Snedecor:1947:Qb

[25] George W. Snedecor. 51. Query. Biometrics, 3(3):137–138, September 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001949.

Baten:1947:Q

[26] W. D. Baten. 52. Query. Biometrics, 3(3):138–139, September 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001950.

Snedecor:1947:Qc

[27] George W. Snedecor. 53. Query. Biometrics, 3(3):139–140, September 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001951.

Anonymous:1947:CQ

[28] Anonymous. Correction for query (46). Biometrics, 3(3):140, September 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001952. See [15].

DeLury:1947:EBP

[29] D. B. DeLury. On the estimation of biological populations. Biometrics, 3 (4):145–167, December 1947. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001390.

Packard:1947:WMB

[30] Charles Packard. Welcome to the marine biological laboratory. Biometrics, 3(4):168–171, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001391.

Rasch:1947:RBD

[31] G. Rasch. Recent biometric developments in Denmark. Biometrics, 3 (4):172–175 + 189, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001392.

Bose:1947:RWI

[32] R. C. Bose. Recent work on “incomplete block designs” in India. Biometrics, 3(4):176–178, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001393.

Belz:1947:BWA

[33] Maurice Belz. Biometric work in Australia. Biometrics, 3(4):179, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001394.

Neurdenburg:1947:BDN

[34] M. G. Neurdenburg. Biometric developments in The Netherlands. Biometrics, 3(4):179, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001395.

Anonymous:1947:D

[35] Anonymous. Discussion. Biometrics, 3(4):180–183, December 1947. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001396.

Teissier:1947:BSC

[36] Georges Teissier. The Biometric Society — constitution. Biometrics, 3 (4):184–187, December 1947. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001397.

Teissier:1947:ACP

[37] George Teissier. Allocution de cloture de la Première Conference Internationale de Biometrie. (French) [Closing remarks of the First International Biometrics Conference]. Biometrics, 3(4):187–189, December 1947. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001398.

Bliss:1947:BSR

[38] C. I. Bliss. The Biometric Society. report of council meetings. Biometrics, 3(4):190–192, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001399.

Anonymous:1947:Ac

[39] Anonymous. Abstracts. Biometrics, 3(4):193, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001400.

Kempthorne:1947:Qb

[40] Oscar Kempthorne. 54. Query. Biometrics, 3(4):194, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001401.

Federer:1947:Q

[41] Walter T. Federer. 55. Query. Biometrics, 3(4):195, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001402.

Anonymous:1947:NN

[42] Anonymous. News and notes. Biometrics, 3(4):196–198, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001403.

Anonymous:1947:VI

[43] Anonymous. Volume information. Biometrics, 3(4):199–200, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001388.

Anonymous:1947:FM

[44] Anonymous. Front matter. Biometrics, 3(4):??, December 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001389.

Fisher:1948:QTG

[45] R. A. Fisher. A quantitative theory of genetic recombination and chiasma formation. Biometrics, 4(1):1–13, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001694. See errata [73].

Teissier:1948:RDS

[46] Georges Teissier. La relation D’Allometrie: Sa signification statistique et biologique. (French) [The allometry relationship: Its statistical and biological significance]. Biometrics, 4(1):14–53, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001695.

Kempthorne:1948:GTPa

[47] Oscar Kempthorne and Walter T. Federer. The general theory of primepower lattice designs. I. Introduction and designs for p^{n} varieties in blocks of p plots. Biometrics, 4(1):54–79, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001696.

Snedecor:1948:Qa

[48] George W. Snedecor. 56. Query. Biometrics, 4(1):80–81, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001697.

Anonymous:1948:BSa

[49] Anonymous. The Biometric Society. Biometrics, 4(1):82–85, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001698.

Anonymous:1948:AMB

[50] Anonymous. Annual meeting of the Biometrics Section. Biometrics, 4(1): 86, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001699.

Anonymous:1948:SSS

[51] Anonymous. Statistical summer sessions at the University of California. Biometrics, 4(1):87–88, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001700.

Anonymous:1948:FMa

[52] Anonymous. Front matter. Biometrics, 4(1):??, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001693.

McCulloch:1948:SON

[53] Warren S. McCulloch and Walter Pitts. The statistical organization of nervous activity. Biometrics, 4(2):91–99, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001453.

Wadley:1948:EDC

[54] F. M. Wadley. Experimental design in comparison of allergens on cattle. Biometrics, 4(2):100–108, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001454.

Kempthorne:1948:GTPb

[55] O. Kempthorne and W. T. Federer. The general theory of prime-power lattice designs. II. Designs for p^{n} varieties in blocks of p^{s} plots, and in squares.

Biometrics, 4(2):109–121, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001455.

Young:1948:AIO

[56] D. M. Young and R. G. Romans. Assays of insulin with one blood sample per rabbit per test day. Biometrics, 4(2):122–131, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001456.

Snedecor:1948:Qb

[57] George W. Snedecor. 57. Query. Biometrics, 4(2):132–134, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001457.

Bancroft:1948:Q

[58] T. A. Bancroft. 58. Query. Biometrics, 4(2):135–136, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001458.

Anonymous:1948:A

[59] Anonymous. Abstracts. Biometrics, 4(2):137–144, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001459.

Anonymous:1948:BSb

[60] Anonymous. The Biometric Society. Biometrics, 4(2):145–147, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001460.

Anonymous:1948:FMb

[61] Anonymous. Front matter. Biometrics, 4(2):??, June 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001452.

DeLury:1948:AC

[62] D. B. DeLury. The analysis of covariance. Biometrics, 4(3):153–170, September 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001562.

Anderson:1948:EDU

[63] R. L. Anderson and H. L. Manning. An experimental design used to estimate the optimum planting date for cotton. Biometrics, 4(3):171–196, September 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001563.

Thompson:1948:UPN

[64] William R. Thompson. On the use of parallel or nonparallel systems of transformed curves in bioassay: Illustration in the quantitative complement-fixation test. Biometrics, 4(3):197–210, September 1948. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001564.

Snedecor:1948:Qc

[65] George W. Snedecor. 59. Query. Biometrics, 4(3):211–213, September 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001565.

Anonymous:1948:Q

[66] Anonymous. 60. Query. Biometrics, 4(3):213–215, September 1948. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001566.

Fisher:1948:BS

[67] R. A. Fisher. The Biometric Society. Biometrics, 4(3):216–219, September 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001567. See errata [73].

Anonymous:1948:NN

[68] Anonymous. News and notes. Biometrics, 4(3):220–222, September 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001568.

Anonymous:1948:FMc

[69] Anonymous. Front matter. Biometrics, 4(3):??, September 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001561.

Nordskog:1948:SRS

[70] A. W. Nordskog and S. Lee Crump. Systematic and random sampling for estimating egg production in poultry. Biometrics, 4(4):223–233, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001407.

Rigney:1948:SAC

[71] J. A. Rigney and R. E. Blaser. Sampling alyce clover for chemical analyses. Biometrics, 4(4):234–239, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001408.

Quenouille:1948:ACN

[72] M. H. Quenouille. The analysis of covariance and non-orthogonal comparisons. Biometrics, 4(4):240–246, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001409.

Anonymous:1948:EQT

[73] Anonymous. Errata: A quantitative theory of genetic recombination and chiasma formation. Biometrics, 4(4):246, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001410. See [45].

Rao:1948:FPC

[74] C. Radhakrishna Rao and D. C. Shaw. On a formula for the prediction of cranial capacity. Biometrics, 4(4):247–253, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001411.

Comstock:1948:CGV

[75] R. E. Comstock and H. F. Robinson. The components of genetic variance in populations of biparental progenies and their use in estimating the average degree of dominance. Biometrics, 4(4):254–266, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001412.

Horvitz:1948:Q

[76] Daniel G. Horvitz. 61. Query. Biometrics, 4(4):267–269, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001413.

Snedecor:1948:Qd

[77] George W. Snedecor. 62. Query. Biometrics, 4(4):269–271, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001414.

Snedecor:1948:Qe

[78] George W. Snedecor. 63. Query. Biometrics, 4(4):271–272, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001415.

Anonymous:1948:BSc

[79] Anonymous. The Biometric Society. Biometrics, 4(4):273–276, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001416.

Anonymous:1948:VI

[80] Anonymous. Volume information. Biometrics, 4(4):281–283, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001405.

Anonymous:1948:FMd

[81] Anonymous. Front matter. Biometrics, 4(4):??, December 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001406.

Harshbarger:1949:TRL

[82] Boyd Harshbarger. Triple rectangular lattices. Biometrics, 5(1):1–13, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001888.

Miller:1949:POC

[83] Richard B. Miller. Problems of the optimum catch in small whitefish lakes. Biometrics, 5(1):14–26, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001889.

Fry:1949:SLT

[84] F. E. J. Fry. Statistics of a lake trout fishery. Biometrics, 5(1):27–67, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001890.

Anderson:1949:Q

[85] R. L. Anderson. 64. Query. Biometrics, 5(1):68–71, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001891.

Ostle:1949:Q

[86] Bernard Ostle. 65. Query. Biometrics, 5(1):71–75, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001892.

Anonymous:1949:A

[87] Anonymous. Abstracts. Biometrics, 5(1):76–89, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001893.

Wood:1949:BS

[88] Eric C. Wood. The Biometric Society. Biometrics, 5(1):90–93, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001894.

Anonymous:1949:FMa

[89] Anonymous. Front matter. Biometrics, 5(1):??, March 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001887.

Tukey:1949:CIM

[90] John W. Tukey. Comparing individual means in the analysis of variance. Biometrics, 5(2):99–114, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/

3001913.

Rounsefell:1949:MET

[91] George A. Rounsefell. Methods of estimating total runs and escapements of salmon. Biometrics, 5(2):115–126, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001914.

Mather:1949:AET

[92] K. Mather. The analysis of extinction time data in bioassay. Biometrics, 5(2):127–143, June 1949. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001915.

Federer:1949:GTP

[93] Walter T. Federer. The general theory of prime-power lattice designs. III. The analysis for p^{3} varieties in blocks of p plots with more than 3 replicates. Biometrics, 5(2):144–161, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001916.

Quenouille:1949:RBL

[94] M. H. Quenouille. A relation between the logarithmic, Poisson, and negative binomial series. Biometrics, 5(2):162–164, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001917.

Anscombe:1949:SAI

[95] F. J. Anscombe. The statistical analysis of insect counts based on the negative binomial distribution. Biometrics, 5(2):165–173, June 1949. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001918.

Anonymous:1949:Qa

[96] Anonymous. 66. Query. Biometrics, 5(2):174, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001919.

Anonymous:1949:Qb

[97] Anonymous. 67. Query. Biometrics, 5(2):175–176, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001920.

Anonymous:1949:Qc

[98] Anonymous. 68. Query. Biometrics, 5(2):176–178, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001921.

Bliss:1949:Q

[99] C. I. Bliss. 69. Query. Biometrics, 5(2):178–179, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001922.

Anonymous:1949:BSa

[100] Anonymous. The Biometric Society. Biometrics, 5(2):180–182, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001923.

Anonymous:1949:FMb

[101] Anonymous. Front matter. Biometrics, 5(2):??, June 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001912.

Marcuse:1949:OAV

[102] Sophie Marcuse. Optimum allocation and variance components in nested sampling with an application to chemical analysis. Biometrics, 5(3):189– 206, September 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL https://www.jstor.org/stable/3001935.

Bartlett:1949:FSL

[103] M. S. Bartlett. Fitting a straight line when both variables are subject to error. Biometrics, 5(3):207–212, September 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001936.

Schuck:1949:RCC

[104] Howard A. Schuck. Relationship of catch to changes in population size of New England haddock. Biometrics, 5(3):213–231, September 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001937.

Tukey:1949:ODF

[105] John W. Tukey. One degree of freedom for non-additivity. Biometrics, 5 (3):232–242, September 1949. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001938.

Chassan:1949:SAI

[106] J. B. Chassan. On a statistical approximation to the infection interval. Biometrics, 5(3):243–249, September 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/3001939.

Anonymous:1949:Qd

[107] Anonymous. 70. Query. Biometrics, 5(3):250–251, September 1949. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001940.

Anonymous:1949:Qe

[108] Anonymous. 71. Query. Biometrics, 5(3):251–252, September 1949. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001941.

Sprague:1949:Q

[109] G. F. Sprague. 72. Query. Biometrics, 5(3):252–253, September 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001942.

Anonymous:1949:BSb

[110] Anonymous. The Biometric Society. Biometrics, 5(3):254–255, September 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001943.

Anonymous:1949:FMc

[111] Anonymous. Front matter. Biometrics, 5(3):??, September 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001934.

Finney:1949:CRM

[112] D. J. Finney. The choice of a response metameter in bio-assay. Biometrics, 5(4):261–272, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 3001511. See correction [122].

Jerne:1949:VMR

[113] N. K. Jerne and Eric C. Wood. The validity and meaning of the results of biological assays. Biometrics, 5(4):273–299, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001512.

Fisher:1949:BAT

[114] R. A. Fisher. A biological assay of tuberculins. Biometrics, 5(4):300–316, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001513.

Masuyama:1949:ODD

[115] Motosaburo Masuyama. On a one-dimensional diffusion method of assaying antibiotic substances and its fundamental formulas. Biometrics, 5(4): 317–329, December 1949. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/3001514.

Healy:1949:RCB

[116] M. J. R. Healy. Routine computation of biological assays involving a quantitative response. Biometrics, 5(4):330–334, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001515.

Finney:1949:Q

[117] D. J. Finney. 73. Query. Biometrics, 5(4):335–337, December 1949. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001516.

Tukey:1949:Q

[118] John W. Tukey. 74. Query. Biometrics, 5(4):337–340, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001517. See correction [123].

Anonymous:1949:BSc

[119] Anonymous. The Biometric Society. Biometrics, 5(4):341–344, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001518.

Anonymous:1949:VI

[120] Anonymous. Volume information. Biometrics, 5(4):349–353, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001509.

Anonymous:1949:FMd

[121] Anonymous. Front matter. Biometrics, 5(4):??, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https:

//www.jstor.org/stable/3001510.

Anonymous:1950:CCR

[122] Anonymous. Correction: The choice of a response metameter in bio-assay. Biometrics, 6(1):33, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/stable/ 3001422. See [112].

Anonymous:1950:CQ

[123] Anonymous. Correction: In query 74. Biometrics, 6(2):168, June 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001499. See [118].